Evaluate the indefinite integral
∫(1x7+85x)dx

+ C
1
Expert's answer
2020-12-03T18:33:24-0500

"\int(17x+85x)dx=\int(102x)dx=\frac{102x^2}{2}+C=\\=51x^2+C, C\in constant"